# Data-Efficient 3D Visual Grounding via Order-Aware Referring

### v3

@@ Appendix E Visualization of Responses in Each Referring Block
Figure A7: Visualization of Responses in Each Referring Block. An example in NR3D is shown in this figure. We color each object proposal according to their feature response in $F_{1:(B+1)}$, where a larger response with a brighter color. Note that the blue bounding box represents the ground truth target object. We can see that our Vigor progressively locates the target object by considering the referential order by first focusing on the wall then the screen and finally the trash can.
To show that our Vigor indeed progressively locates the target object following the derived referential order, we visualize the feature response of $F_{1:(B+1)}$ ($B=4$) in A7. The blue bounding box indicates the ground-truth target object, and we color the object proposal according to the response of their corresponding features in $F_{1:(B+1)}$, where a brighter color represents a higher response. We can see that the responses to object proposals are originally cluttered. As our referential blocks are applied, the response of anchor/target objects becomes larger and finally locates the ideal target object in the last feature $F_{5}$.
@@ Appendix F Details and Prefix Prompt Examples of Two-Stage ICL for Deriving Referential Order
Figure A8: Referential order generation via LLM. A two-stage ICL is deployed to remove redundant information from $D$, forming a descriptive order $O_{1:B}$ for locating the target object $O_{B}$.
To have Object-Referring blocks $\{R_{1},\cdots,R_{B}\}$ in Sec. 3 to locate the target object properly, it is desirable to extract a proper referential order $O_{1:B}$ from the input description $D$. With such a referring path constructed, visual features of the associated objects can be updated for grounding purposes. This is inspired by the idea of Chain-of-Thoughts [41, 26] in LLM, as noted in CoT3DRef [5]. To achieve such a parsing task, we apply GPT-3.5-Turbo [48] as the description parser using in-context learning (ICL) [14], as depicted in Fig. A8.
@@
To tackle the above problem, we conduct a two-stage in-context learning (ICL) scheme to remove redundant information in $D$ before producing $O_{1:B}$, as depicted in Fig. A8.
@@ F.1 Prefix Prompt of First-Stage ICL
Now for the description [DESCRIPTION], give me the summarized description and the target object. Your answer must be in the form "summarized description: target object:"
@@ F.2 Prefix Prompt of Second-Stage ICL
referential order 1: easy chair->water bottle
@@
referential order 4: desk->pencil->desk lamp->backpack
@@
referential order 6: cabinet->clock->bookshelf
@@
referential order 10: wall->door->towel
Now for the description: [DESCRIPTION], give me the anchor objects and the referential order. Your answer must be in the form "referential order, anchor objects:. "
@@ F.3 Examples of Derived Referential Order
To show that our two-stage ICL produces reasonable referential orders, we provide examples and comparisons between ours and CoT3DRef [5]’s parsing results. In particular, we leverage CoT3DRef’s released prompt to query the GPT-3.5-Turbo. We display 4 examples in A7, where CoT3DRef misses an anchor object “bed” in the third example and includes a redundant object “shelves” as an anchor object in the fourth example, while our Vigor produces proper results in both cases. This shows the effectiveness of the two-stage ICL strategy.
Table A7: Examples of LLM-parsed referential order from CoT3DRef [5] and Vigor. Note that the blue text represents the ideal anchor/target objects, and the red text represents the redundant object that should not appear in the referential order. We can see that MVT misses one anchor object in the third example and includes a redundant object in the fourth example, while Vigor predicts proper order for both cases.
Description CoT3DRef [5] Vigor (Ours)
The pillow closest to the foot of the bed. bed -> pillow bed -> pillow
Facing the bed, it’s the large white pillow on the right. The second one from the headboard. bed -> headboard -> pillow bed -> headboard -> pillow
The front pillow on the bed with the laptop. bed -> pillow laptop -> bed -> pillow
The window near the table, not the one near the shelves. table -> shelves -> window table -> window
@@ G.1.1 LLM-parsed Referential Order
Since Vigor utilizes LLMs to generate the referential order from the input description, despite of our introduced pre-training strategy to warm the training process, the correctness of the extracted referential order would affect the training and the performance of Vigor.

### v4

@@ Appendix E Visualization of Responses in Each Referring Block
Figure 7: Visualization of Responses in Each Referring Block. An example in NR3D is shown in this figure. We color each object proposal according to their feature response in $F_{1:(B+1)}$, where a larger response with a brighter color. Note that the blue bounding box represents the ground truth target object. We can see that our Vigor progressively locates the target object by considering the referential order by first focusing on the wall then the screen and finally the trash can.
To show that our Vigor indeed progressively locates the target object following the derived referential order, we visualize the feature response of $F_{1:(B+1)}$ ($B=4$) in Fig. 7. The blue bounding box indicates the ground-truth target object, and we color the object proposal according to the response of their corresponding features in $F_{1:(B+1)}$, where a brighter color represents a higher response. We can see that the responses to object proposals are originally cluttered. As our referential blocks are applied, the response of anchor/target objects becomes larger and finally locates the ideal target object in the last feature $F_{5}$.
@@ Appendix F Details and Prefix Prompt Examples of Two-Stage ICL for Deriving Referential Order
Figure 8: Referential order generation via LLM. A two-stage ICL is deployed to remove redundant information from $D$, forming a descriptive order $O_{1:B}$ for locating the target object $O_{B}$.
To have Object-Referring blocks $\{R_{1},\cdots,R_{B}\}$ in Sec. 3 to locate the target object properly, it is desirable to extract a proper referential order $O_{1:B}$ from the input description $D$. With such a referring path constructed, visual features of the associated objects can be updated for grounding purposes. This is inspired by the idea of Chain-of-Thoughts [45, 27] in LLM, as noted in CoT3DRef [5]. To achieve such a parsing task, we apply GPT-3.5-Turbo [52] as the description parser using in-context learning (ICL) [14], as depicted in Fig. 8.
@@
To tackle the above problem, we conduct a two-stage in-context learning (ICL) scheme to remove redundant information in $D$ before producing $O_{1:B}$, as depicted in Fig. 8.
@@ F.1 Prefix Prompt of First-Stage ICL
Now for the description [DESCRIPTION], give me the summarized description and the target object. Your answer must be in the form ”summarized description: target object:”
@@ F.2 Prefix Prompt of Second-Stage ICL
[…]
@@
[…]
@@
[…]
Table 8: Examples of LLM-parsed referential order from CoT3DRef [5] and Vigor. Note that the blue text represents the ideal anchor/target objects, and the red text represents the redundant object that should not appear in the referential order. We can see that MVT misses one anchor object in the third example and includes a redundant object in the fourth example, while Vigor predicts proper order for both cases.
Description CoT3DRef [5] Vigor (Ours)
The pillow closest to the foot of the bed. bed $\rightarrow$ pillow bed $\rightarrow$ pillow
Facing the bed, it’s the large white pillow on the right. The second one from the headboard. bed $\rightarrow$ headboard $\rightarrow$ pillow bed $\rightarrow$ headboard $\rightarrow$ pillow
The front pillow on the bed with the laptop. bed $\rightarrow$ pillow laptop $\rightarrow$ bed $\rightarrow$ pillow
The window near the table, not the one near the shelves. table $\rightarrow$ shelves $\rightarrow$ window table $\rightarrow$ window
@@
[…]
Now for the description: [DESCRIPTION], give me the anchor objects and the referential order. Your answer must be in the form ”referential order, anchor objects:. ”
@@ F.3 Examples of Derived Referential Order
To show that our two-stage ICL produces reasonable referential orders, we provide examples and comparisons between ours and CoT3DRef [5]’s parsing results. In particular, we leverage CoT3DRef’s released prompt to query the GPT-3.5-Turbo. We display 4 examples in Table 8, where CoT3DRef misses an anchor object “bed” in the third example and includes a redundant object “shelves” as an anchor object in the fourth example, while our Vigor produces proper results in both cases. This shows the effectiveness of the two-stage ICL strategy.
@@ G.1 Limitations
Table 9: Our adopted GPT-3.5-Turbo’s zero-shot accuracy (%) on identifying the class name of the target object in the referring descriptions. Note that since the ground-truth labels and orders of the anchor objects are not available, we are only allowed to check if the target object is correctly placed at the last position in the parsed referential order as an indirect verification of the reliability of the orders.
Method NR3D SR3D
train test train test
GPT-3.5-Turbo 86.9 89.1 96.4 96.1
@@ G.1.1 LLM-parsed Referential Order
Since Vigor utilizes LLMs to generate the referential order from the input description, despite of our introduced pre-training strategy to warm the training process, the correctness of the extracted referential order would affect the training and the performance of Vigor. For example, Table 9 shows the accuracy on the SR3D dataset and NR3D dataset that our GPT-3.5-Turbo-parsed referential orders correctly place the target object at the last position. We can see that for the NR3D dataset, where the relations in the descriptions are much more complicated, the accuracy of the identification of the target objects is $86.9\%$ and $89.1\%$ for the training and testing sets, respectively. The gap between this accuracy and an absolutely reliable prediction (i.e., $100\%$ accuracy), though small, would still affect the training stability and testing accuracy of our visual grounding pipeline. As a result, better usage of the LLM to produce perfect referential order is one of the future research directions to pursue.